# Leveraging Peer, Self, and Teacher Assessments for Generative AI-Enhanced Feedback

### v1

@@ I Introduction
Formative assessment and feedback are widely recognized as fundamental components of effective learning in higher education. Rather than serving merely as a mechanism for evaluation, feedback functions as a dynamic process that fosters learning, reflection, and performance improvement [morris2021formative]. However, this process cannot rely solely on teacher-provided judgments; it requires authentic evaluative experiences in which students actively engage in giving and receiving feedback [sadler1989formative].
A valuable way to engage students in this evaluative process is through self-assessment, which involves reflecting on and judging the quality of one’s own work against explicit criteria. In particular, when students practice criterion-referenced self-assessment, they become more aware of learning goals, better understand teachers’ expectations, and use the process to identify strengths and areas for improvement. Beyond enhancing performance, self-assessment promotes metacognition and self-regulated learning, as students internalize standards of quality and develop a sense of responsibility for their progress [andrade2007student].
Complementary to this, peer assessment invites students to provide constructive feedback on the work of their classmates. This approach not only helps students recognize errors and refine their work but also allows evaluators to apply shared criteria critically, deepening their understanding and avoiding similar mistakes in their own tasks [topping2009peer]. However, peer feedback also presents several well-documented challenges. Students providing feedback often struggle to compose comments that are sufficiently detailed, balanced, and constructive, partly because of limited experience, uncertainty about quality standards, or a lack of confidence in their evaluative abilities. At the same time, feedback recipients may find it difficult to interpret and apply peer comments, particularly when these address superficial features rather than the substance of the work [wei2024incorporating].
To address these challenges, the integration of Generative Artificial Intelligence (GenAI) offers promising opportunities to enhance feedback processes. For instance, [sajadi2024harnessing] has shown that incorporating tools such as ChatGPT into peer review activities can make feedback more actionable, specific, and constructive, while simultaneously filtering out unhelpful or inappropriate comments. In this context, the AICoFe system (“Artificial Intelligence-based Collaborative Feedback System”) [becerra2025enhancing], developed at Universidad Autonoma de Madrid, integrates rubric-based peer and teacher assessments with a Generative AI model (GePeTo [becerra2024generative]) to generate personalized, high-quality feedback. In addition, AICoFe enables students to perform self-assessment, fostering reflection and the development of both technical (hard) and professional (soft) skills.
@@ II-A Rubrics in Peer, Self, and Teacher Assessment
Teacher assessment has traditionally been at the core of educational evaluation, as teachers are primarily responsible for designing, implementing, and interpreting assessment practices that guide student learning [black2018classroom, brown2022past]. Beyond grading, teacher assessment assumes a formative dimension by informing instructional decisions and providing feedback that promotes student reflection and improvement. Within this context, several studies have highlighted the value of rubrics as essential tools for supporting teachers in articulating learning goals, communicating quality criteria, and ensuring consistency and transparency in feedback [pastore2023teacher].
When effectively designed, rubrics can serve as powerful formative instruments that guide teachers’ qualitative judgments rather than simply quantifying performance. This qualitative orientation enables teachers to provide feedback that is descriptive, constructive, and improvement-oriented instead of relying on aggregated scores. Moreover, the formative use of rubrics helps clarify learning expectations, reduce ambiguity in evaluation, and foster meaningful dialogue between teachers and students about quality standards and progress [panadero2020critical, brookhart2018appropriate].
By making learning criteria explicit and fostering a shared understanding of quality, rubrics not only enhance the consistency of teacher assessment but also lay the foundation for more participatory forms of evaluation. When students are familiar with the criteria used by their teachers, they are better equipped to apply these same standards to assess their own work and that of their peers. In this sense, rubrics act as mediating tools that bridge teacher, self-, and peer assessment, promoting a common evaluative language and supporting the development of students’ evaluative judgment and self-regulated learning [panadero2020critical, andrade2007student]. Likewise, rubrics provide clear reference points that help students interpret quality, offer more focused feedback, and reflect critically on their own performance [reinholz2016assessment].
@@ II-B Comparing Peer, Self, and Teacher Assessment
The relationship among peer, self-, and teacher assessment has been widely examined in higher education to explore issues of validity, reliability, and bias. Early meta-analyses established that peer ratings tend to approximate teacher ratings to a moderate degree. For instance, [falchikov2000student] reported an average correlation of $r=0.69$ across 48 studies in higher education. Their findings highlighted that peer marks align more closely with teacher marks when global judgments based on clearly defined criteria are used, as opposed to fragmented, dimension-specific ratings. Similarly, [li2016peer] reported slightly lower but still substantial average correlation of $r=0.63$ between peer and teacher ratings, confirming that peer assessment can produce reasonably valid results, even in online settings. However, their analysis revealed important moderating factors: peer–teacher agreement is stronger in graduate courses, for individual rather than group work, when assessors are non-anonymous, when feedback includes qualitative comments, and when students are involved in developing rating criteria.
Despite these generally positive findings, several studies point to systematic biases in peer evaluation, such as leniency, friendship, or reciprocity effects, which can undermine reliability [magin2001reciprocity, stonewall2018review]. While [magin2001reciprocity] demonstrated that reciprocity effects account for only a small portion of score variance, later studies highlight the persistence of social and demographic biases, such as gender and cultural stereotypes, particularly in collaborative and team-based learning environments [stonewall2018review].
Comparatively, self-assessment research shows a more nuanced pattern. While self-ratings often diverge from teacher marks, especially among less experienced students, they strongly contribute to metacognitive development and self-regulated learning [ross2006reliability, nieminen2025student]. Recent meta-reviews confirm that self-assessment accuracy improves when rubrics are used and when students receive guided reflection opportunities [tang2023comparison].
@@ II-C GenAI-Supported Feedback and Learning Analytics
Recent research in educational technology have emphasized the transformative potential of Generative Artificial Intelligence (GenAI) in supporting feedback and Learning Analytics. For instance, [giannakos2025promise] highlights that large language models (LLMs) can be leveraged to diagnose learners’ challenges, generate personalized feedback, and enhance self-regulated learning processes, although the authors also warn of significant limitations regarding reliability, ethics, and human oversight.
@@
This growing interest in GenAI-supported learning has led to the rapid development of various educational tools that integrate multimodal generative models. Recent studies indicate that such systems can process multiple input types (e.g., text, image, audio, and video) to provide adaptive feedback, automated assessment, and personalized learning experiences [imran2024google]. For example, [mello2025empowering] presented an approach that leverages Large Language Models (LLMs) to transcribe handwritten student essays and automatically score them according to predefined educational criteria. Their study demonstrates how multimodal LLMs can convert handwritten text into digital form and evaluate writing quality, providing feedback while reducing teachers’ workload. Similarly, [xavier2025empowering] evaluated an AI-driven feedback platform designed to assist instructors in assessing open-ended student responses. The system integrates LLMs and learning analytics to generate personalized, tag-based feedback and recommendations, improving both the quality and efficiency of the assessment process. Their findings show that instructors not only produced more detailed feedback with LLM support but also adopted most of the AI-generated suggestions with minimal modification.
@@
Learning Analytics dashboards [verbert2020learning, becerra2023m2lads] have also evolved through the incorporation of generative AI technologies. For example, [yan2024vizchat] introduced VizChat, a prototype designed to augment traditional dashboards with contextualised explanations generated by multimodal generative AI. VizChat enables users to interact with dashboard visualisations through natural language, receiving personalised and context-aware explanations of complex learning data. This approach addresses the cognitive overload often associated with visual analytics by allowing learners and instructors to query visualisations conversationally, enhancing comprehension and reflective practice. Similarly, [becerra2024generative] proposed GePeTo, a generative AI-based personalized guidance tool integrated into dashboards to provide students with adaptive feedback and support. GePeTo leverages anonymized data and indicators to generate customized guidance messages validated for accuracy and tone before being displayed in the dashboards.
@@
However, as [topali2025designing] point out, the growing use of GenAI and AI-based systems in education requires a stronger human-centered and pedagogically grounded perspective to ensure that these technologies align with pedagogical values and ethical principles.
@@ III-A Context and Participants
The study involved 46 undergraduate students (6 female students) in the final year of the Telecommunication Technology and Service Engineering program at Universidad Autonoma de Madrid. They were enrolled in a compulsory course to train students in enhancing communication and oral presentation skills in engineering contexts.
@@
During the course, a peer assessment methodology was implemented using the AICoFe system [becerra2025enhancing]. As illustrated in Figure 1, students had two roles: evaluator (two peers per presentation) and presenter. The course teacher also acted as evaluator, using the same rubric implemented in AICoFe. All presentations were delivered in a face-to-face classroom setting. For each presentation, the two peer evaluators were randomly assigned to ensure fairness and to avoid systematic bias.
@@ III-D Statistical Analyses
To measure inter-rater agreement, we employed several complementary indices. Weighted Cohen’s kappa was used to estimate pairwise agreement, although its known sensitivity to skewed distributions motivated the additional use of Gwet’s AC2 coefficient, which is more robust in such cases [vach2023gwet]. Kendall’s coefficient of concordance (W) was calculated to assess the global agreement across all evaluators.
@@ V-A Agreement Across Evaluators
The results indicate that when comparing evaluators (Table II), the level of agreement between evaluators was substantial according to Gwet’s AC2 coefficients and relatively consistent across all comparisons, with slightly higher values observed between peer evaluators and between peers and self-assessments. In contrast, the lowest agreement was found between the teacher and the presenters’ self-assessments, a pattern commonly reported in peer assessment literature where students tend to evaluate themselves more leniently than teachers do [tang2023comparison, ross2006reliability, kilic2016examination].
It is important to note that weighted Cohen’s $\kappa$ values were considerably lower, which can be explained by the well-known kappa paradox, a phenomenon where high observed agreement coexists with low $\kappa$ values due to prevalence or bias effects [zec2017high]. For this reason, Gwet’s AC2 provides a more robust and interpretable measure of inter-rater agreement in this context.
@@
An important factor that may have contributed to this level of consistency is the use of a detailed rubric within the AICoFe system, which included clearly defined levels of scale. The rubric was accessible to all participants throughout the course and explicitly discussed during training sessions, fostering a shared understanding of the assessment criteria. This common framework likely helped align interpretations across teacher, peers, and self-assessors, reducing subjectivity and enhancing reliability. Moreover, the substantial agreement observed between peers and between peers and the teacher suggests that students were able to apply the evaluation criteria in a manner broadly consistent with the teacher’s judgments. These findings support previous evidence that well-structured rubrics and transparent assessment procedures can enable peer evaluations to approximate teacher assessments in reliability [kilic2016examination, tang2023comparison].
@@ V-B Correlation Across Evaluations
The correlation analyses offer further evidence of the validity and coherence of the evaluation process. As shown in Table IV, the teacher’s rubric-based scores correlated very strongly with her overall grade ($r=0.88$), confirming internal consistency in the teacher’s judgment. Peer assessments were also significantly correlated with the teacher’s overall grade ($r=0.69$), indicating that students were generally able to recognize similar performance patterns. This value is comparable to the average correlation reported in large-scale meta-analyses of peer and teacher assessments, which range from moderate to substantial across diverse educational contexts [falchikov2000student, li2016peer].
In contrast, self-assessments showed a moderate correlation with the teacher’s grade ($r=0.59$), suggesting that students’ evaluative judgments were partially aligned but still affected by calibration and self-perception biases. This pattern is consistent with previous research showing that students often misjudge their own performance levels, a phenomenon sometimes explained by the Dunning–Kruger effect, whereby lower-performing individuals tend to overestimate their abilities while higher-performing individuals underestimate them [kruger1999unskilled, dunning2011dunning]. Similar trends, strong peer–self relationships but weaker alignment with teacher ratings, have also been reported in professional education contexts [mehrdad2012comparative].
@@ V-C Bias Analysis
The comparative analysis revealed that students tended to assign slightly higher scores to themselves and their peers than those assigned by the teacher, a pattern that has been consistently reported in previous studies [oren2018self]. However, the magnitude of this bias was limited, and differences between peer and self-assessments were not statistically significant (Table VI), suggesting that students’ scoring tendencies were relatively homogeneous.
The higher mean scores in peer and self-assessments reflect a general positive scoring bias, where raters systematically overestimate performance [falchikov2000student, patri2002influence]. In our study, the absence of strong systematic deviations between peer and self-assessment scores, combined with the moderate agreement indices (AC2 = 0.63), suggests that bias was more related to overall score inflation than reciprocal favoritism. These findings are consistent with previous evidence indicating that reciprocity or collusion effects in peer assessment are typically negligible [magin2001reciprocity].
@@ V-D Implications for Generative AI-Enhanced Feedback
The findings of this study offer actionable insights for the evolution of generative AI–based feedback systems such as AICoFe [becerra2025enhancing]. The current implementation of AICoFe integrates teacher and peer assessments as inputs for the generative model GePeTo [becerra2024generative], which produces personalized feedback grounded in rubric-based evaluations. This AI-generated feedback is presented through Learning Analytics dashboards to support students’ self-reflection and improvement. To ensure pedagogical validity and ethical oversight, all generative outputs are reviewed by teachers before being shared with students.

### v2

@@ I Introduction
Formative assessment and feedback are widely recognized as fundamental components of effective learning in higher education. Rather than serving merely as a mechanism for evaluation, feedback functions as a dynamic process that fosters learning, reflection, and performance improvement [22]. However, this process cannot rely solely on teacher-provided judgments; it requires authentic evaluative experiences in which students actively engage in giving and receiving feedback [31].
A valuable way to engage students in this evaluative process is through self-assessment, which involves reflecting on and judging the quality of one’s own work against explicit criteria. In particular, when students practice criterion-referenced self-assessment, they become more aware of learning goals, better understand teachers’ expectations, and use the process to identify strengths and areas for improvement. Beyond enhancing performance, self-assessment promotes metacognition and self-regulated learning, as students internalize standards of quality and develop a sense of responsibility for their progress [1].
Complementary to this, peer assessment invites students to provide constructive feedback on the work of their classmates. This approach not only helps students recognize errors and refine their work but also allows evaluators to apply shared criteria critically, deepening their understanding and avoiding similar mistakes in their own tasks [36]. However, peer feedback also presents several well-documented challenges. Students providing feedback often struggle to compose comments that are sufficiently detailed, balanced, and constructive, partly because of limited experience, uncertainty about quality standards, or a lack of confidence in their evaluative abilities. At the same time, feedback recipients may find it difficult to interpret and apply peer comments, particularly when these address superficial features rather than the substance of the work [39].
To address these challenges, the integration of Generative Artificial Intelligence (GenAI) offers promising opportunities to enhance feedback processes. For instance, [32] has shown that incorporating tools such as ChatGPT into peer review activities can make feedback more actionable, specific, and constructive, while simultaneously filtering out unhelpful or inappropriate comments. In this context, the AICoFe system (“Artificial Intelligence-based Collaborative Feedback System”) [5], developed at Universidad Autonoma de Madrid, integrates rubric-based peer and teacher assessments with a Generative AI model (GePeTo [8]) to generate personalized, high-quality feedback. In addition, AICoFe enables students to perform self-assessment, fostering reflection and the development of both technical (hard) and professional (soft) skills.
@@ II-A Rubrics in Peer, Self, and Teacher Assessment
Teacher assessment has traditionally been at the core of educational evaluation, as teachers are primarily responsible for designing, implementing, and interpreting assessment practices that guide student learning [9, 11]. Beyond grading, teacher assessment assumes a formative dimension by informing instructional decisions and providing feedback that promotes student reflection and improvement. Within this context, several studies have highlighted the value of rubrics as essential tools for supporting teachers in articulating learning goals, communicating quality criteria, and ensuring consistency and transparency in feedback [27].
When effectively designed, rubrics can serve as powerful formative instruments that guide teachers’ qualitative judgments rather than simply quantifying performance. This qualitative orientation enables teachers to provide feedback that is descriptive, constructive, and improvement-oriented instead of relying on aggregated scores. Moreover, the formative use of rubrics helps clarify learning expectations, reduce ambiguity in evaluation, and foster meaningful dialogue between teachers and students about quality standards and progress [26, 10].
By making learning criteria explicit and fostering a shared understanding of quality, rubrics not only enhance the consistency of teacher assessment but also lay the foundation for more participatory forms of evaluation. When students are familiar with the criteria used by their teachers, they are better equipped to apply these same standards to assess their own work and that of their peers. In this sense, rubrics act as mediating tools that bridge teacher, self-, and peer assessment, promoting a common evaluative language and supporting the development of students’ evaluative judgment and self-regulated learning [26, 1]. Likewise, rubrics provide clear reference points that help students interpret quality, offer more focused feedback, and reflect critically on their own performance [29].
@@ II-B Comparing Peer, Self, and Teacher Assessment
The relationship among peer, self-, and teacher assessment has been widely examined in higher education to explore issues of validity, reliability, and bias. Early meta-analyses established that peer ratings tend to approximate teacher ratings to a moderate degree. For instance, [13] reported an average correlation of $r=0.69$ across 48 studies in higher education. Their findings highlighted that peer marks align more closely with teacher marks when global judgments based on clearly defined criteria are used, as opposed to fragmented, dimension-specific ratings. Similarly, [18] reported slightly lower but still substantial average correlation of $r=0.63$ between peer and teacher ratings, confirming that peer assessment can produce reasonably valid results, even in online settings. However, their analysis revealed important moderating factors: peer–teacher agreement is stronger in graduate courses, for individual rather than group work, when assessors are non-anonymous, when feedback includes qualitative comments, and when students are involved in developing rating criteria.
Despite these generally positive findings, several studies point to systematic biases in peer evaluation, such as leniency, friendship, or reciprocity effects, which can undermine reliability [19, 33]. While [19] demonstrated that reciprocity effects account for only a small portion of score variance, later studies highlight the persistence of social and demographic biases, such as gender and cultural stereotypes, particularly in collaborative and team-based learning environments [33].
Comparatively, self-assessment research shows a more nuanced pattern. While self-ratings often diverge from teacher marks, especially among less experienced students, they strongly contribute to metacognitive development and self-regulated learning [30, 23]. Recent meta-reviews confirm that self-assessment accuracy improves when rubrics are used and when students receive guided reflection opportunities [34].
@@ II-C GenAI-Supported Feedback and Learning Analytics
Recent research in educational technology have emphasized the transformative potential of Generative Artificial Intelligence (GenAI) in supporting feedback and Learning Analytics. For instance, [14] highlights that large language models (LLMs) can be leveraged to diagnose learners’ challenges, generate personalized feedback, and enhance self-regulated learning processes, although the authors also warn of significant limitations regarding reliability, ethics, and human oversight.
@@
This growing interest in GenAI-supported learning has led to the rapid development of various educational tools that integrate multimodal generative models. Recent studies indicate that such systems can process multiple input types (e.g., text, image, audio, and video) to provide adaptive feedback, automated assessment, and personalized learning experiences [15]. For example, [21] presented an approach that leverages Large Language Models (LLMs) to transcribe handwritten student essays and automatically score them according to predefined educational criteria. Their study demonstrates how multimodal LLMs can convert handwritten text into digital form and evaluate writing quality, providing feedback while reducing teachers’ workload. Similarly, [40] evaluated an AI-driven feedback platform designed to assist instructors in assessing open-ended student responses. The system integrates LLMs and learning analytics to generate personalized, tag-based feedback and recommendations, improving both the quality and efficiency of the assessment process. Their findings show that instructors not only produced more detailed feedback with LLM support but also adopted most of the AI-generated suggestions with minimal modification.
@@
Learning Analytics dashboards [38, 6] have also evolved through the incorporation of generative AI technologies. For example, [41] introduced VizChat, a prototype designed to augment traditional dashboards with contextualised explanations generated by multimodal generative AI. VizChat enables users to interact with dashboard visualisations through natural language, receiving personalised and context-aware explanations of complex learning data. This approach addresses the cognitive overload often associated with visual analytics by allowing learners and instructors to query visualisations conversationally, enhancing comprehension and reflective practice. Similarly, [8] proposed GePeTo, a generative AI-based personalized guidance tool integrated into dashboards to provide students with adaptive feedback and support. GePeTo leverages anonymized data and indicators to generate customized guidance messages validated for accuracy and tone before being displayed in the dashboards.
@@
However, as [35] point out, the growing use of GenAI and AI-based systems in education requires a stronger human-centered and pedagogically grounded perspective to ensure that these technologies align with pedagogical values and ethical principles.
@@ III-A Context and Participants
The study [3] involved 46 undergraduate students (6 female students) in the final year of the Telecommunication Technology and Service Engineering program at Universidad Autonoma de Madrid. They were enrolled in a compulsory course to train students in enhancing communication and oral presentation skills in engineering contexts.
@@
During the course, a peer assessment methodology was implemented using the AICoFe system [5]. As illustrated in Figure 1, students had two roles: evaluator (two peers per presentation) and presenter. The course teacher also acted as evaluator, using the same rubric implemented in AICoFe. All presentations were delivered in a face-to-face classroom setting. For each presentation, the two peer evaluators were randomly assigned to ensure fairness and to avoid systematic bias.
@@ III-D Statistical Analyses
To measure inter-rater agreement, we employed several complementary indices. Weighted Cohen’s kappa was used to estimate pairwise agreement, although its known sensitivity to skewed distributions motivated the additional use of Gwet’s AC2 coefficient, which is more robust in such cases [37]. Kendall’s coefficient of concordance (W) was calculated to assess the global agreement across all evaluators.
@@ V-A Agreement Across Evaluators
The results indicate that when comparing evaluators (Table II), the level of agreement between evaluators was substantial according to Gwet’s AC2 coefficients and relatively consistent across all comparisons, with slightly higher values observed between peer evaluators and between peers and self-assessments. In contrast, the lowest agreement was found between the teacher and the presenters’ self-assessments, a pattern commonly reported in peer assessment literature where students tend to evaluate themselves more leniently than teachers do [34, 30, 16].
It is important to note that weighted Cohen’s $\kappa$ values were considerably lower, which can be explained by the well-known kappa paradox, a phenomenon where high observed agreement coexists with low $\kappa$ values due to prevalence or bias effects [42]. For this reason, Gwet’s AC2 provides a more robust and interpretable measure of inter-rater agreement in this context.
@@
An important factor that may have contributed to this level of consistency is the use of a detailed rubric within the AICoFe system, which included clearly defined levels of scale. The rubric was accessible to all participants throughout the course and explicitly discussed during training sessions, fostering a shared understanding of the assessment criteria. This common framework likely helped align interpretations across teacher, peers, and self-assessors, reducing subjectivity and enhancing reliability. Moreover, the substantial agreement observed between peers and between peers and the teacher suggests that students were able to apply the evaluation criteria in a manner broadly consistent with the teacher’s judgments. These findings support previous evidence that well-structured rubrics and transparent assessment procedures can enable peer evaluations to approximate teacher assessments in reliability [16, 34].
@@ V-B Correlation Across Evaluations
The correlation analyses offer further evidence of the validity and coherence of the evaluation process. As shown in Table IV, the teacher’s rubric-based scores correlated very strongly with her overall grade ($r=0.88$), confirming internal consistency in the teacher’s judgment. Peer assessments were also significantly correlated with the teacher’s overall grade ($r=0.69$), indicating that students were generally able to recognize similar performance patterns. This value is comparable to the average correlation reported in large-scale meta-analyses of peer and teacher assessments, which range from moderate to substantial across diverse educational contexts [13, 18].
In contrast, self-assessments showed a moderate correlation with the teacher’s grade ($r=0.59$), suggesting that students’ evaluative judgments were partially aligned but still affected by calibration and self-perception biases. This pattern is consistent with previous research showing that students often misjudge their own performance levels, a phenomenon sometimes explained by the Dunning–Kruger effect, whereby lower-performing individuals tend to overestimate their abilities while higher-performing individuals underestimate them [17, 12]. Similar trends, strong peer–self relationships but weaker alignment with teacher ratings, have also been reported in professional education contexts [20].
@@ V-C Bias Analysis
The comparative analysis revealed that students tended to assign slightly higher scores to themselves and their peers than those assigned by the teacher, a pattern that has been consistently reported in previous studies [24]. However, the magnitude of this bias was limited, and differences between peer and self-assessments were not statistically significant (Table VI), suggesting that students’ scoring tendencies were relatively homogeneous.
The higher mean scores in peer and self-assessments reflect a general positive scoring bias, where raters systematically overestimate performance [13, 28]. In our study, the absence of strong systematic deviations between peer and self-assessment scores, combined with the moderate agreement indices (AC2 = 0.63), suggests that bias was more related to overall score inflation than reciprocal favoritism. These findings are consistent with previous evidence indicating that reciprocity or collusion effects in peer assessment are typically negligible [19].
@@ V-D Implications for Generative AI-Enhanced Feedback
The findings of this study offer actionable insights for the evolution of generative AI–based feedback systems such as AICoFe [5]. The current implementation of AICoFe integrates teacher and peer assessments as inputs for the generative model GePeTo [8], which produces personalized feedback grounded in rubric-based evaluations. This AI-generated feedback is presented through Learning Analytics dashboards to support students’ self-reflection and improvement. To ensure pedagogical validity and ethical oversight, all generative outputs are reviewed by teachers before being shared with students.